# Deriving a GENERIC system from a Hamiltonian system

### v2

@@ 5 Conclusion and discussion
While the coarse-graining of Hamiltonian systems is a classical topic, the introduction of the compression hypothesis from dilation theory gives a new type of insight into this question. Specifically, we have shown how the compression hypothesis gives an abstract characterization of heat-bath behaviour that automatically generates two important effects.
@@
In addition, the same compression hypothesis has a clear connection to the GENERIC structure. The operator ${\mathbb{D}}$ that appears in the compression hypothesis generates terms in both the Poisson operator ${\mathbb{J}}_{\mathrm{GEN}}$ and the Onsager operator ${\mathbb{K}}_{\mathrm{GEN}}$: the skew-symmetic part of ${\mathbb{D}}$ contributes to ${\mathbb{J}}_{\mathrm{GEN}}$, and the symmetric part to ${\mathbb{K}}_{\mathrm{GEN}}$. In this way the compression hypothesis generates new structure in the (formerly) Hamiltonian system, which turns out to be GENERIC structure.
@@ 5.1 Relation with models and results in the literature
Coupled Hamiltonian systems consisting of a fixed system and a ‘heat bath’ have been studied by many authors. The earliest example that we could find is by Zwanzig [Zwa73a], and similar systems return in [Zwa05, Zwa01b, KS02b]. In all of these works the heat bath is finite-dimensional; infinite-dimensional heat baths were used in the related work of Jaksic and Pillet [JaP98b] and Rey-Bellet [Rey06].
The fact that a particular tuning of the properties of the heat bath and its projector can lead to Markovianity of the projected subsystem was observed by Ford, Kac, and Mazur [FKM65], and later used by many others [Zwa73b, Zwa05]. Stuart and co-workers [StW99, KS02b] appear to be the first to give a rigorous convergence statement in the limit of infinite dimensions of the heat bath (and the Compression Property 1.3 is effectively equivalent to the choice of the frequencies of the $n$ oscillators in these two papers).
@@
The setup of Jaksic and Pillet [JaP98a] is similar to ours, but their condition (H3) effectively excludes the case of this paper. This is due to the already-remarked fact that ${\mathbf{W}}\cap\mathrm{dom}({\mathbb{J}}_{\mathcal{B}})=\{0\}$, which is proved in Remark LABEL:re:Dom.JB.bfW. Jaksic and Pillet find from their (H3) that their equivalent of the process $Y_{t}$ has $C^{1}$ dependence on the initial datum $(z_{0},\eta_{0})$ (their Theorem 2.1), while because of the lack of continuity we only obtain measurability.
The coarse-graining of this paper is closely related to the ‘Mori-Zwanzig’ reduction method [Mor65, Zwa61, Zwa01a, CHK00, Sch22]. The first step in this method is the expression of the ‘hidden’ variables in terms of the ‘visible’ variables. In in this paper equation (2.10) plays the same role. The integral term characterizes the changes in $z$ that are generated by $z$ itself but through the heat bath:
@@
The Compression Hypothesis (1.5) amounts to assuming that there exists a finite-dimensional subspace ${\mathbf{W}}$ with corresponding orthogonal projection ${\mathbb{P}}\colon{\mathbf{H}}\to{\mathbf{W}}$ and a dissipative operator ${\mathbb{D}}$ on ${\mathbf{W}}$ such that
@@ 5.4 Other derivations of GENERIC by coarse-graining
The closest in spirit to this paper is the coarse-graining approach by Öttinger [Ö05, Ch. 6]. The starting point is again a Hamiltonian system, and a central role is played by a ‘coarse-graining map’ similar to $\pi_{z,w,e}$ in (1.7). The two approaches differ in the definition of the coarse-grained state: while in this paper we consider $\pi_{z,w,e}(z,\eta)$ itself to be the coarse-grained state (up to renormalization of the infinite energy), in [Ö05, Ch. 6] the coarse-grained state is in effect a distribution $\rho_{z,w,e}$ on ${\mathbf{Z}}{\times}{\mathbf{X}}$, parametrized by $(z,w,e)$.
@@ A.1 Basic setup
This generalises the finite-dimensional case, where one can take ${\mathbf{H}}={\mathbf{X}}$ and $\operatorname{cov}_{\mu}=\beta^{-1}\mathrm{id}$, and the properties above reduce to those of Lemma A.1.
@@ B.1 Section 2: Microscopic setup
    As a first step we construct the Gaussian measure $\overline{\gamma}_{\beta}$ on an extension $\overline{\mathbf{X}}$ of the special space $\overline{\mathbf{H}}={\mathrm{L}}^{2}({\mathbb{R}};{\mathbf{W}})$. This leads to a classical white noise process, but now taking values in the finite-dimensional Hilbert space ${\mathbf{W}}$. While translations ${\mathbb{S}}_{t}$ can easily defined for the white noise process, we have to be more careful and construct a Hilbert space $\overline{\mathbf{X}}$ on which we can extend the unitary group $\big{(}\overline{\mathbf{H}},({\mathbb{S}}_{t})_{t\in{\mathbb{R}}}\big{)}$ to a strongly continuous group $\big{(}\overline{\mathbf{X}},({\mathbb{S}}^{\mathbf{X}}_{t})_{t\in{\mathbb{R}}}\big{)}$, which will be no longer unitary. If ${\mathbb{N}}\colon\overline{\mathbf{H}}\to\overline{\mathbf{X}}$ is the embedding operator, then the extension property means

### v3

@@ 5 Conclusion and discussion
While the coarse-graining of Hamiltonian systems is a classical topic, the introduction of the compression property from dilation theory gives a new type of insight into this question. Specifically, we have shown how the compression property gives an abstract characterization of heat-bath behaviour that automatically generates two important effects.
@@
In addition, the same compression property has a clear connection to the GENERIC structure. The operator ${\mathbb{D}}$ that appears in the compression property generates terms in both the Poisson operator ${\mathbb{J}}_{\mathrm{GEN}}$ and the Onsager operator ${\mathbb{K}}_{\mathrm{GEN}}$: the skew-symmetic part of ${\mathbb{D}}$ contributes to ${\mathbb{J}}_{\mathrm{GEN}}$, and the symmetric part to ${\mathbb{K}}_{\mathrm{GEN}}$. In this way the compression property generates new structure in the (formerly) Hamiltonian system, which turns out to be GENERIC structure.
@@ 5.1 Relation with models and results in the literature
Coupled Hamiltonian systems consisting of a fixed system and a ‘heat bath’ have been studied by many authors. The earliest example that we could find is by Zwanzig [Zwa73], and similar systems return in [Zwa05, Zwa01, KS02]. The system of this paper is a generalization of that of [Zwa73] to an infinite-dimensional heat bath. Infinite-dimensional heat baths were also used in the related work of Jaksic and Pillet [JaP98] and Rey-Bellet [Rey06].
The fact that a particular tuning of the properties of the heat bath and its projector can lead to Markovianity of the projected subsystem was observed by Ford, Kac, and Mazur [FKM65], and later used by many others [Zwa73, Zwa05]. Stuart and co-workers [StW99, KS02] appear to be the first to give a rigorous convergence statement in the limit of infinite dimensions of the heat bath (and the Compression Property 1.3 is effectively equivalent to the choice of the frequencies of the $n$ oscillators in these two papers).
@@
The setup of Jaksic and Pillet [JaP98] is similar to ours, but their condition (H3) effectively excludes the case of this paper. This is due to the already-remarked fact that ${\mathbf{W}}\cap\mathrm{dom}({\mathbb{J}}_{\mathcal{B}})=\{0\}$, which is proved in Remark LABEL:re:Dom.JB.bfW. Jaksic and Pillet find from their (H3) that their equivalent of the process $Y_{t}$ has $C^{1}$ dependence on the initial datum $(z_{0},\eta_{0})$ (their Theorem 2.1), while because of the lack of continuity we only obtain measurability.
The coarse-graining of this paper is closely related to the ‘Mori-Zwanzig’ reduction method [Mor65, Zwa61, Zwa01, CHK00, Sch22]. The first step in this method is the expression of the ‘hidden’ variables in terms of the ‘visible’ variables. In this paper equation (2.10) plays the same role. The integral term characterizes the changes in $z$ that are generated by $z$ itself but through the heat bath:
@@
The Compression Property (1.5) amounts to assuming that there exists a finite-dimensional subspace ${\mathbf{W}}$ with corresponding orthogonal projection ${\mathbb{P}}\colon{\mathbf{H}}\to{\mathbf{W}}$ and a dissipative operator ${\mathbb{D}}$ on ${\mathbf{W}}$ such that
@@ 5.4 Other derivations of GENERIC by coarse-graining
The closest in spirit to this paper is the coarse-graining approach by Öttinger [Ött05, Ch. 6]. The starting point is again a Hamiltonian system, and a central role is played by a ‘coarse-graining map’ similar to $\pi_{z,w,e}$ in (1.7). The two approaches differ in the definition of the coarse-grained state: while in this paper we consider $\pi_{z,w,e}(z,\eta)$ itself to be the coarse-grained state (up to renormalization of the infinite energy), in [Ött05, Ch. 6] the coarse-grained state is in effect a distribution $\rho_{z,w,e}$ on ${\mathbf{Z}}{\times}{\mathbf{X}}$, parametrized by $(z,w,e)$.
@@ A.1 Basic setup
This generalizes the finite-dimensional case, where one can take ${\mathbf{H}}={\mathbf{X}}$ and $\operatorname{cov}_{\mu}=\beta^{-1}\mathrm{id}$, and the properties above reduce to those of Lemma A.1.
@@ B.1 Section 2: Microscopic setup
    As a first step we construct the Gaussian measure $\overline{\gamma}_{\beta}$ on an extension $\overline{\mathbf{X}}$ of the special space $\overline{\mathbf{H}}={\mathrm{L}}^{2}({\mathbb{R}};{\mathbf{W}})$. This leads to a classical white noise process, but now taking values in the finite-dimensional Hilbert space ${\mathbf{W}}$. While translations ${\mathbb{S}}_{t}$ can easily be defined for the white noise process, we have to be more careful and construct a Hilbert space $\overline{\mathbf{X}}$ on which we can extend the unitary group $\big{(}\overline{\mathbf{H}},({\mathbb{S}}_{t})_{t\in{\mathbb{R}}}\big{)}$ to a strongly continuous group $\big{(}\overline{\mathbf{X}},({\mathbb{S}}^{\mathbf{X}}_{t})_{t\in{\mathbb{R}}}\big{)}$, which will be no longer unitary. If ${\mathbb{N}}\colon\overline{\mathbf{H}}\to\overline{\mathbf{X}}$ is the embedding operator, then the extension property means